# MetaVLA: Unified Meta Co-Training for Efficient Embodied Adaptation

### v1

@@ 1 Introduction
Building on these ideas, we propose MetaVLA, a unified framework that fills a critical gap in VLA post-training by intelligently introducing auxiliary tasks without incurring the inefficiencies of per-task SFT or the performance drop of naive multi-task SFT. It introduces Context-Aware Meta Co-Training, a streamlined paradigm that optimizes target tasks while infusing extrinsic signals from auxiliary tasks to elevate performance. MetaVLA integrates VLA with a meta-learning co-training strategy to enhance efficiency and generalization in limited-data adaptation. Specifically, a single model is trained across target tasks (e.g., LIBERO suites) to harness cross-task gradients, while auxiliary tasks are incorporated through a memory-augmented mechanism derived from Attentive Neural Processes (ANP) Kim et al. (2019). This lightweight module injects out-of-domain information gain without disrupting target optimization, enabling scalable and robust adaptation. MetaVLA is maintenance-friendly, backbone-agnostic, and easily extends beyond SFT to training paradigms like reinforcement learning. Figure 1 highlights three key advantages of MetaVLA over existing approaches.
@@ 2.2 Multi-Task Co-training
Co-training across tasks has long been used to improve generalization Doersch & Zisserman (2017); Zhang & Yang (2021), scalability Devlin et al. (2019); Sun et al. (2020); McLean et al. (2025), and data efficiency Aghajanyan et al. (2021); Crawshaw (2020). More recently, it has shown strong success in LLMs and VLMs. GPT-2 Radford et al. (2019), for example, leverages diverse pretraining sources (e.g., web pages, Wikipedia, news) for broad generalization. LLaVA Liu et al. (2023b), a pioneering open-source VLM, uses multitask fine-tuning for multimodal alignment across conversation, captioning, and reasoning tasks. This trend continues in models like Qwen-3 Yang et al. (2025), which expands co-training diversity by incorporating code, textbooks, and multilingual data across both pretraining and post-training. Similarly, Molmo and Pixmo Deitke et al. (2024) provide detailed ablations on co-training with varied data sources, demonstrating the benefits of task and domain diversity. These advances highlight co-training as a key driver of performance in both pretraining and post-training stages.
@@ 2.3 Meta-Learning
Meta-learning enables models to quickly adapt to new tasks, often using diverse contextual data and through episodic training Finn et al. (2017); Koch (2015); Santoro et al. (2016); Ravi & Larochelle (2016). Attentive Neural Processes (ANP) Kim et al. (2019), an amortized meta-learners inspired by Gaussian Processes, learn a distribution over functions conditioned on both global prior and target-specific latent vectors via attention mechanism Vaswani et al. (2023). ANP is well-suited for VLA due to its task-invariance, selective attention to relevant demonstrations, and avoidance of direct context optimization during adaptation. These properties simplify cross-domain training, enhance stability, and enable scalability—crucial for leveraging auxiliary data effectively, as shown in later results.
Figure 2: MetaVLA Architecture. VLA backbone married with Context-Aware Meta Co-Training Framework, where the context memory bank is composed of both in-domain target tasks and out-of-domain auxiliary tasks.
@@ 3.2.1 Architecture
Unlike standard ANP, which uses smaller-scale neural networks, we integrate a pretrained Llama-2 Touvron et al. (2023) backbone from OpenVLA. MAR generates both stochastic and deterministic contextual latent vectors, which are concatenated with the Llama hidden states before the final output layer. The combined representations are then passed through the LM head to produce output logits, enabling end-to-end training via standard Llama decoding. See Figure 2 for an overview of the framework.
@@ 3.3 Auxiliary Tasks Selection
LIBERO tasks feature a Franka Emika Panda arm with a gripper and primarily use front-facing camera views. In contrast, selected GR00T tasks include bimanual manipulation using front views and single-arm manipulation with side views only. These variations are intentionally chosen to test the robustness and generalization ability of MetaVLA. An example of task difference among these tree types is showing in Figure 3, and more examples are in Section A.2.
@@ 4.1 Experiment setting
Table 1: Success rate comparison with prior methods. All MetaVLA variants are single models trained for 75K steps. MetaVLA (ours) uses only LIBERO suites in the context bank without the stochastic module, while MetaVLA+Stochastic (ours) includes it. SFT-4LIBERO is a single-model baseline trained with vanilla multi-task SFT across all suites. OpenVLA (top) comprises four Hugging Face models fine-tuned separately on LIBERO using the OpenVLA-7B backbone, totaling roughly 240K steps. MetaVLA with six auxiliary tasks surpasses OpenVLA by 4.4% and SFT-4LIBERO by 3.1% on average, with even larger gains on LIBERO-Long (8.0% and 5.1%, respectively).
@@ A.4.2 Effect of Context Batch Size
A.5 Success Cases in LIBERO Simulation
Figures 10, 11, 12, and 13 demonstrate example execution sequences of MetaVLA successfully completing one task from each LIBERO suite in its simulation: Goal, Spatial, Object, and Long.

### v3

@@ 1 Introduction
Building on these ideas, we propose MetaVLA, a unified framework that fills a critical gap in VLA post-training by intelligently introducing auxiliary tasks without incurring the inefficiencies of per-task SFT or the performance drop of naive multi-task SFT. It introduces Context-Aware Meta Co-Training, which jointly trains all target tasks with a unified model, improving adaptation by leveraging cross-task data through a context bank. The context bank is a memory-augmented mechanism that contains auxiliary knowledge with domain diversities derived from Attentive Neural Processes (ANP) (Kim et al., 2019) based on Meta-learning. This lightweight module injects out-of-domain information gain without disrupting target optimization, enabling scalable and robust adaptation. MetaVLA is maintenance-friendly, backbone-agnostic, and easily extends beyond SFT to training paradigms like reinforcement learning. Figure 1 highlights three key advantages of MetaVLA over existing approaches.
@@ 2.2 Multi-Task Co-training
Co-training across tasks has long been used to improve generalization Doersch and Zisserman (2017); Zhang and Yang (2021), scalability Devlin et al. (2019); Sun et al. (2020); McLean et al. (2025), and data efficiency Aghajanyan et al. (2021); Crawshaw (2020). More recently, it has shown strong success in LLMs and VLMs. GPT-2 Radford et al. (2019), for example, leverages diverse pretraining sources (e.g., web pages, Wikipedia, news) for broad generalization. LLaVA Liu et al. (2023b), a pioneering open-source VLM, uses multitask fine-tuning for multimodal alignment across conversation, captioning, and reasoning tasks. This trend continues in models like Qwen-3 Yang et al. (2025), which expands co-training diversity by incorporating code, textbooks, and multilingual data across both pretraining and post-training. Similarly, Molmo and Pixmo Deitke et al. (2024) provide detailed ablations on co-training with varied data sources, demonstrating the benefits of task and domain diversity. These advances highlight co-training as a key driver of performance in both pretraining and post-training stages.
@@ 2.3 Meta-Learning
Meta-learning enables models to quickly adapt to new tasks, often using diverse contextual data and through episodic training Finn et al. (2017); Koch (2015); Santoro et al. (2016); Ravi and Larochelle (2016). Attentive Neural Processes (ANP) Kim et al. (2019), an amortized meta-learners inspired by Gaussian Processes, learn a distribution over functions conditioned on both global prior and target-specific latent vectors via attention mechanism Vaswani et al. (2023). ANP is well-suited for VLA due to its task-invariance, selective attention to relevant demonstrations, and avoidance of direct context optimization during adaptation. These properties simplify cross-domain training, enhance stability, and enable scalability—crucial for leveraging auxiliary data effectively, as shown in later results.
Figure 2: MetaVLA Architecture. VLA backbone married with Context-Aware Meta Co-Training Framework, where the context memory bank is composed of both in-domain target tasks and out-of-domain auxiliary tasks. We further detail the definitions of variables in Section 3.2.1.
@@ 3.2 MetaVLA
MetaVLA employs a Context-Aware Meta Co-Training approach that jointly trains on all in-domain suites with a single model, while leveraging contextual demonstrations through meta-learning. We formalize this mechanism by showing how MAR offers a principled way to aggregate and condition on heterogeneous context data, using it to enhance stability during co-training across task suites.
@@ 3.2.1 Architecture
Unlike standard ANP, which uses smaller-scale neural networks, we integrate a pretrained Llama-2 Touvron et al. (2023) backbone from OpenVLA. MAR generates both stochastic and deterministic contextual latent vectors, which are concatenated with the Llama hidden states before the final output layer. The combined representations are then passed through the LM head to produce output logits, enabling end-to-end training via standard Llama decoding. See Figure 2 for an overview of the framework. We further summarize the symbols and definitions in Table 7.
@@ 3.3 Auxiliary Tasks Selection
LIBERO tasks feature a Franka Emika Panda arm with a gripper and primarily use front-facing camera views. In contrast, selected GR00T tasks include bimanual manipulation using front views and single-arm manipulation with side views only. These variations are intentionally chosen to test the robustness and generalization ability of MetaVLA. An example of task difference among these three types is showing in Figure 3, and more examples are in Section A.2.
@@ 4.1 Experiment setting
Table 1: Success rate comparison with prior methods. All MetaVLA variants are single models trained for 75K steps. MetaVLA (ours) uses only LIBERO suites in the context bank without the stochastic module, while MetaVLA+Stochastic (ours) includes it. Method+NSingle+Mbimanual includes NN single arm and $M$ bimanual (two arms) auxiliary tasks described in Section 3.3. SFT-4LIBERO is a single-model baseline trained with vanilla multi-task SFT across all suites. OpenVLA (top) comprises four Hugging Face models fine-tuned separately on LIBERO using the OpenVLA-7B backbone, totaling roughly 240K steps. MetaVLA with six auxiliary tasks surpasses OpenVLA by 4.4% and SFT-4LIBERO by 3.1% on average, with even larger gains on LIBERO-Long (8.0% and 5.1%, respectively).
@@ 4.4 Ablation Study
4.4.1 Effect of Different Backbone
Model Goal Spatial Object Long Average
NORA-Long (Hung et al., 2025) 85.4 90.5 95.0 70.6 85.4
NORA-Long-SFT-4LIBERO 87.0 92.5 94.0 75.5 87.3
NORA-Long-SFT-4LIBERO+5single+1bimanual 73.6 79.5 75.2 37.2 66.4
MetaVLA-NORA-Long (ours) 90.8 96.2 96.5 77.8 90.3
MetaVLA-NORA-Long+5single+1bimanual (ours) 93.8 95.8 97.2 80.2 91.8
Table 2: Success rate comparison of applying MetaVLA variants to NORA-Long. NORA-Long-SFT-4LIBERO is a single-model baseline trained with vanilla multi-task SFT across all suites using NORA-Long. NORA-Long comprises four Hugging Face models fine-tuned separately on LIBERO using the NORA-Long backbone. Without auxiliary tasks, MetaVLA-NORA-Long surpasses NORA-Long by 4.9%, and exceeds NORA-Long-SFT-4LIBERO by 3.0%. Adding auxiliary tasks further pushes the average success rate to 91.8%, achieving the highest performance in LIBERO-Goal, Object, and Long.
To validate MetaVLA’s effectiveness in various backbones, we assessed MetaVLA variants on NORA (Hung et al., 2025), a 3B Qwen2.5-VL-based (Bai et al., 2025) VLA model. We selected NORA-Long (Deep Cognition and Language Research (DeCLaRe) Lab, 2025) variant as the base model because it provides a stronger LIBERO performance baseline than NORA. To ensure fair comparison, we re-evaluate the NORA-Long baselines in the LIBERO simulation environment using the four single-task fine-tuned models from Hugging Face (Deep Cognition and Language Research (DeCLaRe) Lab, 2025), and adopt these as our baselines.
As shown in Table 2, without auxiliary data, MetaVLA outperforms NORA-Long by 4.9% on average and NORA-Long-SFT-4LIBERO by 3.0%. When auxiliary tasks are incorporated, the average success rate further improves by 6.4% compared to NORA-Long. For more challenging suites—Goal and Long—the improvements are 8.4% and 9.6%, respectively. Moreover, consistent with results using the OpenVLA backbone, MetaVLA-NORA-Long+5single+1bimanual significantly outperforms its native SFT counterpart, NORA-Long-SFT-4LIBERO+5single+1bimanual, by 25.4%. The training convergence curves on accuracy and loss in Figure 10 further bolster the stronger stability of MetaVLA during training as more diverse tasks are added. Together, these results demonstrate MetaVLA’s backbone-agnostic capability.
@@ A.4.2 Effect of Context Batch Size
A.4.3 Effect of Different Backbone
Figure 10 shows Training Accuracy and Imitation Loss for 5 single and 1 bimanual auxiliary tasks when using NORA-Long (Hung et al., 2025) backbone. MetaVLA consistently converges to higher performance across both metrics compared to vanilla SFT, proving that our observation on OpenVLA (OpenVLA Contributors, 2024) is also true under another backbone.
Figure 10: Training convergence comparison for NORA-Long backbone with and without MetaVLA after adding 5 single and 1 bimanual auxiliary tasks. Training Accuracy and Imitation Loss are compared between MetaVLA variants and baseline SFT-4LIBERO under 5single+1bimanual co-training settings. Without MetaVLA, vanilla SFT with auxiliary tasks fails to adapt effectively, while proposed MetaVLA consistently achieves better accuracy and lower loss throughout the training.
A.5 Symbols and Definitions
We summarize all the symbols used in our MetaVLA architecture in Table 7
Symbol Name Definition
$\mathbf{x}_{T}$ Target feature Encoded input feature for target queries.
$\mathbf{y}_{T}$ Target action Action corresponding to target query $\mathbf{x}_{T}$.
$\mathbf{x}_{C}$ Context features Encoded input features for context queries.
$\mathbf{y}_{C}$ Context actions Action corresponding to context query $\mathbf{x}_{C}$.
$(x_{Ci},y_{Ci})$ Context pair $i$ A single feature–action pair from the context bank.
$\mathbf{r}_{Ci}$ Deterministic context rep. Self-attention rep. for context pair i.
$\mathbf{r}_{T}$ Deterministic target rep. Cross-attention rep. of $x_{T}$ attending to $\{x_{Ci}\}$.
$\mathbf{s}_{Ci}$ Stochastic context rep. Self-attention rep. used to compute latent posterior.
$\mathbf{\bar{s}}_{C}$ Mean stochastic context rep. Averaged stochastic rep. over all context rep. $\{\mathbf{s}_{Ci}\}$.
$\mathbf{\bar{s}}_{T}$ Mean stochastic target rep. Averaged stochastic rep. over all target rep. $\{\mathbf{s}_{Ti}\}$.
$z$ Latent variable Stochastic latent sampled from approximate posterior.
$q(z|\mathbf{\bar{s}}_{C})$ Context posterior Approx. posterior over $z$ conditioned on context.
$q(z|\mathbf{\bar{s}}_{T})$ Target posterior Training posterior used for variational objective.
$p(\mathbf{y}_{T}|\mathbf{x}_{T},\mathbf{r}_{T},z)$ Decoder likelihood Conditional distribution predicting target actions.
$D_{\mathrm{KL}}(\cdot\|\cdot)$ KL divergence Regularizer aligning target and context posteriors.
Table 7: Symbol table for MAR and MetaVLA. rep. is abbreviation for representation.
A.6 Success Cases in LIBERO Simulation
Figures 11, 12, 13, and 14 demonstrate example execution sequences of MetaVLA successfully completing one task from each LIBERO suite in its simulation: Goal, Spatial, Object, and Long.